OK so the equation is balanced and is:
3Mg + 2AlCl(3) ---> 3MgCl(2) + 2Al
I am given:
mass of crucible and Mg: 34.61g
mass of crucible alone: 28.43g
mass of crucible and Al: 32.42g

I was told to subtract the crucible weight from the other 2 measurements so Mg = 6.18g and Al = 3.99g

IThe questions are:
1) What mass of Mg is used in the reaction?
2) What is the mass of Al produced?
3) What is the Theoretical Yield of AL?
4) What is the percent yield of the reaction?
1
Expert's answer
2012-06-22T08:38:43-0400
1) m(Mg) = m(crucible and Mg) - m(crucible alone) = 34.61 - 28.43 = 6.18 g
2)
m(Al) = m(crucible and Al) - m(crucible alone) = 32.42 - 28.43 = 3.99 g
3)
m(Al, theor.) = m(Mg)*2*M(Al) / 3*M(Mg) = 6.18*2*27 / (24*3) = 4.635 g
4) the
percent yield is (3.99 / 4.635)*100% = 86.1%